Determine if the statement is TRUE or FALSE. Justify your answer. All numbers under discussion are

integers.


5. For each m ≥ 1 and n ≥ 1, if mn is a multiple of 4, then m or n is a multiple of 4.

6. For each m ≥ 1 and n ≥ 1, if mn is a multiple of 3, then m or n is a multiple of 3


1
Expert's answer
2021-02-24T06:01:41-0500

5) FALSE

Let "m=2" and "n=2" .

Number "mn=2\cdot 2=4" is a multiple of "4" , but "m" and "n" aren’t multiples of "4" .

6) TRUE

Suppose to the contrary that "m" and "n" are not multiples of 3. 

This leads to three cases:

1."\ m=3k+1,\ n=3l+1"

"mn=(3k+1)(3l+1)=9kl+3k+3l+1=3(3kl+k+l)+1"

The remainder on dividing "mn" by "3" equals "1" .


2."\ m=3k+1,\ n=3l+2" "mn=(3k+1)(3l+2)=9kl+6k+3l+2=3(3kl+2k+l)+2"

The remainder on dividing "mn" by "3" equals "2" .


3."\ m=3k+2,\ n=3l+2" "mn=(3k+2)(3l+2)=9kl+6k+6l+1=3(3kl+2k+2l+1)+1"

The remainder on dividing "mn" by "3" equals "1" .


In each of cases, we have that "mn" is not a multiple of 3. It follows that if "m" and "n" are not multiples of "3", then "mn" is not a multiple of "3".


Therefore,  if "mn" is a multiple of "3" , then "m" or "n"  is a multiple of "3" .